# Learning Beyond Experience: Generalizing to Unseen State Space with Reservoir Computing

### v1

@@ Abstract
A plethora of machine learning (ML) techniques have shown impressive success in modeling complex dynamical systems from observed data alone – a challenge of broad practical importance, e.g., in climate science, public health, economics, ecology, and neuroscience. However, these so-called ‘black-box’ models, which do not incorporate guidance based on known or suspected features of the system, often require extensive training data that cover the range of possible system behaviors. When the training domain is incomplete, they often fail outside of it, whereas models that incorporate knowledge of the system’s properties may still succeed. Here, we show that reservoir computing – a black-box ML approach commonly used to model dynamical systems – can overcome this limitation in important settings. By studying systems that can evolve toward multiple distinct long-term behaviors, we show that reservoir computers can make accurate predictions about behaviors they’ve never seen before, even when trained on limited data.
@@ I Introduction
In this paper, we describe a scheme to train RCs on a collection of disjoint time series, allowing for more flexible and exhaustive use of available data. This multiple-trajectory training has previously been applied to multi-task learning, where the goal is to train a single RC across multiple dynamical systems, each of which exhibit different dynamics.Norton et al. (2025); Kong et al. (2021); Panahi and Lai (2024); Kong, Brewer, and Lai (2024); Kim et al. (2020); Lu and Bassett (2020) Here, we leverage the scheme’s flexibility to improve sampling of the state space in multistable dynamical systems with short-lived transients. Then, we utilize the multistability of these systems to test when RCs can generalize from their training data to capture system dynamics in unexplored regions of state space. Specifically, we show that an RC trained on trajectories from a single basin of attraction can recover the dynamics in other unseen basins, capturing even fractal-like basin structures.
@@ II Challenges in Predicting Multistable Systems
where ${\bm{u}(t)}$ is a driving signal. To train an RC for forecasting, we drive the reservoir system with an observed time series that is a function of the state of the true system, ${\bm{u}(t)=\bm{g}(\bm{x}(t))}$, in the ‘open-loop mode’ (Fig. 1). Then we choose a linear readout matrix or ‘output layer’, $W_{out}$, such that $\bm{u}(t)$ can be approximated through a linear projection of the auxiliary state:
@@
Once we have a suitable output layer, the RC can evolve as an autonomous dynamical system, using its own output as the driving signal in the ‘closed-loop mode’ (Fig. 1), to mimic the system of interest:
@@ III Training a Reservoir Computer on Multiple Trajectories
where the $\tanh()$ function is applied element-wise. The input weight matrix, $W_{in}$, couples the $N_{in}$-dimensional input ${\bm{u}(t)}$ to the reservoir nodes. The directed and weighted adjacency matrix, $W_{r}$, specifies the strength and sign of interactions between each pair of nodes, and a random vector of biases, $\bm{b}$, breaks symmetries in the nodes’ dynamics. We say that the reservoir has ‘memory’ if its state ${\bm{r}(t+\Delta t)}$ depends not only on the most recent input, ${\bm{u}(t)}$, but also (recursively) on previous inputs, ${\bm{u}(t-m\Delta t)}$ for ${m>0}$ – i.e., if ${(1-\lambda)}$ and/or the matrix $W_{r}$ are nonzero. The leakage rate, $\lambda$, thus influences the time-scale on which the reservoir state evolves, and, consequently, the duration of its memory.
@@
To train an RC for a forecasting task, we choose its output layer, of dimension ${N_{in}\times N_{r}}$, so that at every time step over a set of ${N_{train}}$ training signals, ${\bm{u}^{(i)},\,i=1,...,N_{train}}$, which we standardize such that each component has mean zero and range one (as measured across the union of all training signals), the RC’s output closely matches its input at the next time step:
@@
where ${\mathrm{RMS}[u_{j}]}$ is the root-mean-square amplitude of the $j^{\mathrm{th}}$ component of the inputs calculated over all training time series, ${\mathcal{N}\left(0,\xi\right)}$ draws a random sample from a Gaussian distribution with mean zero and standard deviation $\xi$, and $\eta$ is a small constant – the ‘noise amplitude.’ We then drive the reservoir with the noisy training signals in the open-loop mode (Eqs. 2, 5 and 1) and minimize the ridge-regression cost function:
@@
[…]
@@
Once the RC has been trained as described above, we use it to obtain predictions, $\hat{\bm{u}}(t)$, of the system of interest. During the prediction phase, the RC operates in closed-loop mode: at each time step, its input is set to its own output from the previous step, allowing it to evolve as an autonomous dynamical system. Used in this way, the RC is intended to mimic the behavior of the system of interest as in Eq. 4:
@@
In typical applications, we wish to predict how the system of interest will evolve, having observed its recent behavior over some period. In this case, we naturally use these recent observations to initialize the forecast. Namely, we drive the reservoir in open-loop mode (Eq. 5) with a short ‘test signal,’ $\bm{u}_{test}$, consisting of the available recent observations and then switch to the closed-loop mode (Eq. 10b) to forecast from the end of $\bm{u}_{test}$. The test signal enables the state of the auxiliary reservoir system to synchronize to the state of the underlying system of interest. In general, an appropriate test signal can be substantially shorter than would be sufficient to train an RC accurately. Hence, an RC that has been trained to accurately capture the dynamics of $\bm{u}(t)$, can be used to predict from a different initial condition by starting from a comparatively short test signal. The test signal should, however, be at least as long as the RC’s memory to ensure that the memory is appropriately initialized. (A few recent studies have also proposed methods to ‘cold-start’ forecasts with test signals that are even shorter than the RC’s memory.Norton et al. (2025); Grigoryeva et al. (2024))
@@ IV Results
We evaluate the ability of our RC setup to generalize to unexplored regions of state space using simulated data from two multistable dynamical systems: the Duffing systemDuffing (1918) and the magnetic pendulum system.Motter et al. (2013) Each of these systems is low-dimensional and dissipative, exhibiting only fixed-point attractors (rather than, e.g., limit cycles or strange attractors). Nonetheless, previous studiesGöring et al. (2024); Zhang and Cornelius (2023) have shown that learning the dynamics of even these simple multistable systems remains challenging for typical RC approaches.
To evaluate the performance of our RC implementation in this context, we adopt a working definition for when a time series $\bm{u}(t)$ is said to converge to a fixed point $\bm{A}^{i}$, ${\bm{A}(\bm{u})=\bm{A}^{i}}$. Specifically, if $\bm{A}^{i}$ is the nearest stable fixed point to the final point of the series, $\bm{u}(t_{f})$, our convergence criteria require that $\bm{u}(t)$ satisfies one of two additional conditions, depending on whether the system state is fully or partially measured. When $\bm{u}(t)$ contains full system state information at every time step, we require that the energy of the system at $t_{f}$, $E\left[\bm{u}(t_{f})\right]$, is below the potential barrier, $E_{0}$, between the system’s stable attractors: ${E\left[\bm{u}(t_{f})\right]<E_{0}}$. When the full system state is not available and we cannot calculate the system energy, we instead require that the final $25$ points of ${\bm{u}(t)}$ are all within a threshold distance, $\varepsilon_{c}$, of the attracting fixed point $\bm{A}^{i}$.
@@
IV.1 Duffing System
Figure 2: RC identification of an unseen attractor in the Duffing system. (a) We train an RC with ${N_{r}=75}$ nodes on ${N_{train}=10}$ fully-observed trajectories (gray lines) from one of the Duffing system’s basins of attraction. (b) Then, we forecast the system’s evolution from $36$ initial conditions (dots) and corresponding short test signals (thick lines, ${N_{test}=10}$ observations). Predictions illustrated by thin lines. We color each trajectory according the true basin of its initial conditions. The RC recovers system behavior in both the seen (blue) and unseen (pink) basins. Only one of the sample predictions (a pink trajectory) goes to the incorrect fixed point.
The unforced Duffing systemDuffing (1918) models an oscillator moving under a nonlinear elastic force with linear damping. It is governed by a pair of coupled ordinary differential equations:
@@ IV.1 Duffing System
With ${a=-1/2}$, ${b=-1}$, and ${c=1/10}$, the Duffing system is dissipative and multistable, with two attracting fixed points,
$\bm{A}^{\pm}=\left(\pm\sqrt{-b/c},0\right)\approx\left(\pm 3.16,0\right),$ (14)
@@
and an unstable fixed point at the origin. Trajectories starting from almost any initial condition ${(x_{0},y_{0})}$ will thus converge to one of the attractors, $\bm{A}^{\pm}$. The initial conditions from which trajectories converge to $\bm{A}^{+}$ form the basin of attraction ${B(\bm{A}^{+})}$ and the initial conditions from which trajectories converge to $\bm{A}^{-}$ form the basin of attraction ${B(\bm{A}^{-})}$. Only trajectories that move along the stable manifold of the unstable fixed point do not converge to one of the two attractors. The corresponding set of initial conditions, which is of measure zero, forms the boundary between ${B(\bm{A}^{+})}$ and ${B(\bm{A}^{-})}$.
@@
Figure 3: RCs can generalize to unseen basins even with sparse and restricted training data. We train an RC of ${N_{r}=75}$ nodes on ${N_{train}=10}$ partially-observed trajectories ($x$ only) of the Duffing system from the basin of attraction $B(\bm{A}^{-})$. Then we make predictions from short test signals (${N_{test}=10}$ observations each) from both basins of attraction. We draw the initial conditions, $(x_{0},y_{0})$, of the training signals from a random distribution that uniformly covers the intersection of $B(\bm{A}^{-})$ with a square of side $2\Delta_{0}^{train}$ centered at the origin (dashed boxes), i.e. ${x_{0}\sim\mathcal{U}\left[-\Delta_{0}^{train},\Delta_{0}^{train}\right]}$ and ${y_{0}\sim\mathcal{U}\left[-\Delta_{0}^{train},\Delta_{0}^{train}\right]}$, subject to $(x_{0},y_{0})\in B(\bm{A}^{-})$. (a) and (b): Example predictions for $\Delta_{0}^{train}=10$. The test signal ends at the vertical dashed line. (c) to (f): The RC-predicted basins for different values of $\Delta_{0}^{train}$, from ${\Delta_{0}^{train}=10}$ (c) to ${\Delta_{0}^{train}=4}$ (f). The blue/pink initial conditions are those that the RC correctly predicts belong to the seen/unseen basin. The initial conditions of test signals for which the RC-predicted trajectory converges to the incorrect attractor are yellow. White initial conditions indicate that the predicted trajectory converges to a spurious attractor. Crosses mark the true fixed point attractors and dots mark the training initial conditions. We use $y_{0}$ for plotting purposes only; the RC has access to and then predicts only the $x$-component of the Duffing system.
@@
We show in Fig. 2 that a reservoir computer (RC) trained on trajectories from only one basin of attraction, ${B(\bm{A}^{-})}$, can capture the Duffing system’s dynamics in both basins of attraction and infer the existence of the other fixed point attractor, which is unseen in the training data. Specifically, we train the RC on ${N_{train}=10}$ different trajectories, each of which converges to the fixed point $\bm{A}^{-}$ (Fig. 2a). We then evaluate the trained RC on short test signals with initial conditions sampled from both basins of attraction (Fig. 2b). Each test signal consists of the first ${N_{test}=10}$ observations of the true system’s evolution from a new initial condition, representing the kind of partial trajectory typically available in a prediction task. These test signals serve both to initialize the forecast and to allow the reservoir to synchronize to the new system state before entering the autonomous closed-loop mode. Even though the RC’s training data explores only the basin of attraction of the fixed point $\bm{A}^{-}$, the RC infers the existence and location of the second attractor, $\bm{A}^{+}$, and correctly predicts that $\bm{A}^{+}$ is also a fixed point (and not, for example, a limit cycle).
@@
In Fig. 3a-c, we train an RC on the same $N_{train}=10$ trajectories from the basin of attraction $B(\bm{A}^{-})$ as shown in Fig. 2, but give the RC access to just the $x$-component of each trajectory, so that it receives only partial information of the system state at every time step. Then, we again forecast from test signals in both basins of attraction, each consisting of $N_{test}=10$ observations of the true system’s evolution. The sample forecasts in Fig. 3a and b highlight the challenging nature of the basin prediction problem. The trajectories of the true system in both cases are almost identical until the system approaches the unstable fixed point at the origin around time step $500$. In Fig. 3a, the RC correctly predicts that the trajectory converges to the unseen fixed point, $\bm{A}^{+}$. In Fig. 3b, however, the true trajectory converges to the seen fixed point, $\bm{A}^{-}$, while the RC predicts that it converges to the unseen fixed point. As the system approaches the stable manifold of the unstable fixed point at the origin (in this case, it approaches the origin itself), the seemingly small prediction error is sufficient to push the autonomous RC system into the incorrect basin of attraction, and the predicted and true trajectories then separate rapidly.
@@
In Fig. 3c we make predictions from test signals with initial conditions ${(x_{0},y_{0})}$ arranged in a ${150\times 150}$ grid spanning ${-10\leq x_{0}\leq 10}$ and ${-10\leq y_{0}\leq 10}$, and plot the RC-predicted basin structure. As before, the RC has access to just the $x$-component of each test signal; we use $y_{0}$ for plotting purposes only. Test signals for which the RC correctly predicts that the system converges to the seen and the unseen attractors are colored blue and pink. Yellow indicates test signals for which the RC predicts the incorrect attractor. The ${N_{train}=10}$ initial conditions of the training signals, all from the basin ${B(\bm{A}^{-})}$, are marked by black dots and the positions of the two attracting fixed points are marked by black crosses.
@@
Fig. 3c demonstrates clearly that an RC trained on data from only one basin of attraction is able to generalize to the unexplored basin. The RC not only infers the existence of the unseen fixed point, but it also achieves high accuracy in predicting whether any given initial condition belongs to the basin $B(\bm{A}^{-})$ or the basin $B(\bm{A}^{+})$. Only near the basin boundary, where a trajectory’s final state is most sensitive to perturbations, does the RC struggle to make accurate basin predictions.
@@
Figure 4: In many scenarios, RC performance in unseen basins is comparable to performance in the training basin. We train an RC of ${N_{r}=75}$ nodes to forecast partially-observed states of the Duffing system ($x$ only) and plot the fraction of short test signals for which the RC predicts the correct basin of attraction as we vary the number of training initial conditions, $N_{train}$, and the half-widths of the ranges of the training and test initial conditions, ${\Delta_{0}^{train}}$ and ${\Delta_{0}^{test}}$. In (a) and (c), all training signals are from the basin of attraction ${B(\bm{A}^{-})}$, and in (b) and (d), training signals are from both basins, ${B(\bm{A}^{\pm})}$. In all cases, we randomly sample the initial conditions of the training trajectories, but the initial conditions of the test signals, which belong to both basins, form a ${50\times 50}$ uniform grid. Each test signal consists of the first ${N_{test}=10}$ consecutive data points from the true system, starting from the specified initial condition. For all grid points, we plot the mean fraction correct calculated over ten random realizations of the RC’s internal structure and of the initial conditions of the training signals. (a) and (b) We fix the number of training initial conditions, ${N_{train}=10}$. (c) and (d) We fix the half-width of the test initial condition range, ${\Delta_{0}^{test}=10}$. Vertical lines mark the distance of the fixed points from the origin, $||\bm{A}^{\pm}||$, and diagonal lines mark ${\Delta_{0}^{test}=\Delta_{0}^{train}}$. Crosses and asterisks indicate, respectively, the values of $N_{train}$, $\Delta_{0}^{train}$, and $\Delta_{0}^{test}$ that we use in Fig. 3a-c and in Fig. 3d-f.

### v2

@@ Abstract
A plethora of machine learning (ML) techniques have shown impressive success in modeling complex dynamical systems from observed data alone – a challenge of broad practical importance, e.g., in climate science, public health, economics, ecology, and neuroscience. However, these so-called ‘black-box’ models, which do not incorporate guidance based on known or suspected features of the system, often require extensive training data that cover the range of possible system behaviors. When the training domain is incomplete, they often fail outside of it, whereas models that incorporate knowledge of the system’s properties may still succeed. Here, we show that reservoir computing – a black-box ML approach commonly used to model dynamical systems – can overcome this limitation in important settings. By studying systems that can evolve toward multiple distinct long-term behaviors, we show that reservoir computers can make accurate predictions about behaviors they have never seen before, even when trained on limited data.
@@ I Introduction
In this paper, we describe a scheme to train RCs on a collection of disjoint time series, allowing for more flexible and exhaustive use of available data. This multiple-trajectory training has previously been applied to multi-task learning, where the goal is to train a single RC across multiple dynamical systems, each of which exhibit different dynamics.Norton et al. (2025); Kong et al. (2021); Panahi and Lai (2024); Kong, Brewer, and Lai (2024); Kim et al. (2020); Lu and Bassett (2020) Here, we leverage the scheme’s flexibility to improve sampling of the state space in multistable dynamical systems with short-lived transients. Then, we utilize the multistability of these systems to test when RCs can generalize from their training data to capture system dynamics in unexplored regions of state space. Specifically, we show that an RC trained on trajectories from a single basin of attraction can recover the dynamics in other unexplored basins, capturing even fractal-like basin structures and unseen chaotic attractors.
@@ II Challenges in Predicting Multistable Systems
where $\bm{u}(t)$} is a driving signal. To train an RC for forecasting, we drive the reservoir system with an observed time series that is a function of the state of the true system, $\bm{u}(t)=\bm{g}(\bm{x}(t))$}, in the ‘open-loop mode’ (Fig.˜1). Then we choose a linear readout matrix or ‘output layer’, $W_{out}$, such that $\bm{u}(t)$ can be approximated through a linear projection of the auxiliary state:
@@
Once we have a suitable output layer, the RC can evolve as an autonomous dynamical system, using its own output as the driving signal in the ‘closed-loop mode’ (Fig.˜1), to mimic the system of interest:
@@ III Training a Reservoir Computer on Multiple Trajectories
where the $\tanh()$ function is applied element-wise. The input weight matrix, $W_{in}$, couples the $N_{in}$-dimensional input $\bm{u}(t)$} to the reservoir nodes. The directed and weighted adjacency matrix, $W_{r}$, specifies the strength and sign of interactions between each pair of nodes, and a random vector of biases, $\bm{b}$, breaks symmetries in the nodes’ dynamics. We say that the reservoir has ‘memory’ if its state $\bm{r}(t+\Delta t)$} depends not only on the most recent input, $\bm{u}(t)$}, but also (recursively) on previous inputs, $\bm{u}(t-m\Delta t)$} for m>0{m>0} – i.e., if $(1-\lambda)$} and/or the matrix $W_{r}$ are nonzero. The leakage rate, $\lambda$, thus influences the time-scale on which the reservoir state evolves, and, consequently, the duration of its memory.
@@
To train an RC for a forecasting task, we choose its output layer, of dimension $N_{in}\times N_{r}$}, so that at every time step over a set of $N_{train}$} training signals, $\bm{u}^{(i)},\,i=1,...,N_{train}$}, which we standardize such that each component has mean zero and maximum absolute value one (as measured across the union of all training signals), the RC’s output closely matches its input at the next time step:
@@
where $\mathrm{RMS}[u_{j}]$} is the root-mean-square amplitude of the $j^{\mathrm{th}}$ component of the inputs calculated over all training time series, $\mathcal{N}\left(0,\xi\right)$} draws a random sample from a Gaussian distribution with mean zero and standard deviation $\xi$, and $\eta$ is a small constant – the ‘noise amplitude.’ We then drive the reservoir with the noisy training signals in the open-loop mode (Eqs.˜2, 5 and 1) and minimize the ridge-regression cost function:
@@
where $N_{i}$ is the number of (evenly-spaced) data points in the $i^{\mathrm{th}}$ signal (i.e., it has duration $(N_{i}-1)\Delta t$}), the scalar $\alpha$ is a (TikhonovTikhonov et al. (1995)) regularization parameter which prevents over-fitting, $\|\ \|$ denotes the Euclidean ($L^{2}$) norm, and $N_{fit}=\sum_{i=1}^{N_{train}}N_{i}-N_{train}(N_{trans}+1)$} is the number of input/output pairs used for fitting. Importantly, we discard the first $N_{trans}$ reservoir states and target outputs of each training signal as a transient to allow the reservoir state to synchronize to each signal before fitting over the remaining time steps. (Note that Eq.˜8 reduces to the usual cost function for single-trajectory training when $N_{train}=1$}.) The minimization problem, Eq.˜8, has solution
@@
Once the RC has been trained as described above, we use it to obtain predictions, $\hat{\bm{u}}(t)$, of the system of interest. During the prediction phase, the RC operates in closed-loop mode: at each time step, its input is set to its own output from the previous step, allowing it to evolve as an autonomous dynamical system. Used in this way, the RC is intended to mimic the behavior of the system of interest as in Eq.˜4:
@@
In typical applications, we wish to predict how the system of interest will evolve, having observed its recent behavior over some period. In this case, we naturally use these recent observations to initialize the forecast. Namely, we drive the reservoir in open-loop mode (Eq.˜5) with a short ‘test signal,’ $\bm{u}_{test}$, consisting of the available recent observations and then switch to the closed-loop mode (Eq.˜10b) to forecast from the end of $\bm{u}_{test}$. The test signal enables the state of the auxiliary reservoir system to synchronize to the state of the underlying system of interest. In general, an appropriate test signal can be substantially shorter than would be sufficient to train an RC accurately. Hence, an RC that has been trained to accurately capture the dynamics of $\bm{u}(t)$, can be used to predict from a different initial condition by starting from a comparatively short test signal. The test signal should, however, be at least as long as the RC’s memory to ensure that the memory is appropriately initialized. (A few recent studies have also proposed methods to ‘cold-start’ forecasts with test signals that are even shorter than the RC’s memory.Norton et al. (2025); Grigoryeva et al. (2024))
@@ IV Results
We evaluate the ability of our RC setup to generalize to unexplored regions of state space using simulated data from four multistable dynamical systems: the Duffing system,Duffing (1918) a multi-well system with segregated basins, a magnetic pendulum system with fractal-like basins,Motter et al. (2013) and a multistable Lorenz-like system with coexisting chaotic attractors.Lü, Chen, and Cheng (2004) Previous studiesGöring et al. (2024); Zhang and Cornelius (2023) have shown that learning the dynamics of even low-dimensional multistable systems can be challenging for typical RC approaches.
To evaluate the performance of our RC implementation in the context of dissipative systems with fixed-point attractors, we adopt a working definition for when a time series $\bm{u}(t)$ is said to converge to a fixed point $\bm{A}^{i}$, $\bm{A}(\bm{u})=\bm{A}^{i}$}. Specifically, if $\bm{A}^{i}$ is the nearest stable fixed point to the final point of the series, $\bm{u}(t_{f})$, our convergence criteria require that $\bm{u}(t)$ satisfies one of two additional conditions, depending on whether the system state is fully or partially measured. When $\bm{u}(t)$ contains full system state information at every time step, we require that the energy of the system at $t_{f}$, $E\left[\bm{u}(t_{f})\right]$, is below the potential barrier, $E_{0}$, between the system’s stable attractors: $E\left[\bm{u}(t_{f})\right]<E_{0}$}. When the full system state is not available and we cannot calculate the system energy, we instead require that the final $25$ points of $\bm{u}(t)$} are all within a threshold distance, $\varepsilon_{c}$, of the attracting fixed point $\bm{A}^{i}$.
@@
IV.1 Intertwined Basins in the Duffing System
Figure 2: RC identification of an unseen attractor in the Duffing system. (a) We train an RC with $N_{r}=200$} nodes on $N_{train}=10$} fully-observed trajectories (gray lines) from one of the Duffing system’s basins of attraction. (b) Then, we forecast the system’s evolution from $36$ initial conditions (dots) and corresponding short test signals (thick lines, $N_{test}=10$} observations). Predictions are illustrated by thin lines. We color each trajectory according the true basin of its initial condition. The RC recovers system behavior in both the seen (blue) and unseen (pink) basins, and predicts the correct fixed point for all sample test signals.
The Duffing systemDuffing (1918) models an oscillator moving under a nonlinear elastic force with linear damping. It is governed by a pair of coupled ordinary differential equations:
@@ IV.1 Intertwined Basins in the Duffing System
With a=1/2{a=-1/2}, b=1{b=-1}, and c=1/10{c=1/10}, the Duffing system is dissipative and multistable, with two attracting fixed points. When the system is unforced ($F_{0}=0$), the positions of these attractors are
$\bm{A}^{\pm}=\left(\pm\sqrt{-b/c},0\right)\approx\left(\pm 3.16,0\right).$ (14)
@@
There is also an unstable fixed point at the origin. Trajectories starting from almost any initial condition $(x_{0},y_{0})$} will thus converge to one of the attractors, $\bm{A}^{\pm}$. The initial conditions from which trajectories converge to $\bm{A}^{+}$ form the basin of attraction $B(\bm{A}^{+})$} and the initial conditions from which trajectories converge to $\bm{A}^{-}$ form the basin of attraction $B(\bm{A}^{-})$}. Only trajectories that move along the stable manifold of the unstable fixed point do not converge to one of the two attractors. The corresponding set of initial conditions, which is of measure zero, forms the boundary between $B(\bm{A}^{+})$} and $B(\bm{A}^{-})$}.
@@
Figure 3: RCs can generalize to unseen basins even with sparse and restricted training data. (a) to (e) We train an RC of $N_{r}=200$} nodes on $N_{train}=10$} partially-observed trajectories ($x$ only) of the unforced Duffing system ($F_{0}=0$) from the basin of attraction $B(\bm{A}^{-})$ (blue). Then we make predictions from short test signals ($N_{test}=10$} observations each) from both basins of attraction. We draw the initial conditions, $(x_{0},y_{0})$, of the training signals from a random distribution that uniformly covers the intersection of $B(\bm{A}^{-})$ with a square of side $2\Delta_{0}^{train}$ centered at the origin (dashed boxes), i.e. $x_{0}\sim\mathcal{U}\left[-\Delta_{0}^{train},\Delta_{0}^{train}\right]$} and $y_{0}\sim\mathcal{U}\left[-\Delta_{0}^{train},\Delta_{0}^{train}\right]$}, subject to $(x_{0},y_{0})\in B(\bm{A}^{-})$. (a) to (c): The RC-predicted basins for different values of $\Delta_{0}^{train}$, from $\Delta_{0}^{train}=10$} (a) to $\Delta_{0}^{train}=4$} (c). The blue/pink initial conditions are those that the RC correctly predicts belong to the seen/unseen basin. The initial conditions of test signals for which the RC-predicted trajectory converges to the incorrect attractor are yellow. White initial conditions indicate that the predicted trajectory converges to a spurious attractor. Crosses mark the true fixed point attractors and dots mark the training initial conditions. We use $y_{0}$ for plotting purposes only; the RC has access to and then predicts only the $x$-component of the Duffing system. (d) and (e): Example predictions for $\Delta_{0}^{train}=10$. The test signal ends at the vertical dashed line. (f) and (g) The RC still generalizes to unseen basins when a constant external forcing ($F_{0}=1$}) breaks the Duffing system’s rotational symmetry, whether we train the RC on the smaller basin (f) or the larger basin (g).
@@
We show in Fig.˜2 that a reservoir computer (RC) trained on trajectories from only one basin of attraction, $B(\bm{A}^{-})$}, can capture the unforced Duffing system’s dynamics in both basins of attraction and infer the existence of the other fixed-point attractor, which is unseen in the training data. Specifically, we train the RC on $N_{train}=10$} different trajectories, each of which converges to the fixed point $\bm{A}^{-}$ (Fig.˜2a). We then evaluate the trained RC on short test signals with initial conditions sampled from both basins of attraction (Fig.˜2b). Each test signal consists of the first $N_{test}=10$} observations of the true system’s evolution from a new initial condition, representing the kind of partial trajectory typically available in a prediction task. These test signals serve both to initialize the forecast and to allow the reservoir to synchronize to the new system state before entering the autonomous closed-loop mode. Even though the RC’s training data explores only the basin of attraction of the fixed point $\bm{A}^{-}$, the RC infers the existence and location of the second attractor, $\bm{A}^{+}$, and correctly predicts that $\bm{A}^{+}$ is also a fixed point (and not, for example, a limit cycle).
@@
In Fig.˜3a-e, we likewise train an RC on $N_{train}=10$ trajectories from the basin of attraction $B(\bm{A}^{-})$, but give the RC access to just the $x$-component of each trajectory, so that it receives only partial information of the system state at every time step. Then, we again forecast from test signals in both basins of attraction, each consisting of $N_{test}=10$ observations of the true system’s evolution. The sample forecasts in Fig.˜3d-e highlight the challenging nature of the basin prediction problem. The trajectories of the true system in both cases are almost identical until the system approaches the unstable fixed point at the origin around time step $500$. In Fig.˜3d, the RC correctly predicts that the trajectory converges to the unseen fixed point, $\bm{A}^{+}$. In Fig.˜3e, however, the true trajectory converges to the seen fixed point, $\bm{A}^{-}$, while the RC predicts that it converges to the unseen fixed point. As the system approaches the stable manifold of the unstable fixed point at the origin (in this case, it approaches the origin itself), the seemingly small prediction error is sufficient to push the autonomous RC system into the incorrect basin of attraction, and the predicted and true trajectories then separate rapidly.
@@
In Fig.˜3a we make predictions from test signals with initial conditions $(x_{0},y_{0})$} arranged in a $150\times 150$} grid spanning $10\leq x_{0}\leq 10$} and $10\leq y_{0}\leq 10$}, and plot the RC-predicted basin structure. As before, the RC has access to just the $x$-component of each test signal; we use $y_{0}$ for plotting purposes only. Test signals for which the RC correctly predicts that the system converges to the seen and the unseen attractors are colored blue and pink. Yellow indicates test signals for which the RC predicts the incorrect attractor. The $N_{train}=10$} initial conditions of the training signals, all from the basin $B(\bm{A}^{-})$}, are marked by black dots and the positions of the two attracting fixed points are marked by black crosses.
@@
Fig.˜3a demonstrates clearly that an RC trained on data from only one basin of attraction is able to generalize to the unexplored basin. The RC not only infers the existence of the unseen fixed point, but it also achieves high accuracy in predicting whether any given initial condition belongs to the basin $B(\bm{A}^{-})$ or the basin $B(\bm{A}^{+})$. Only near the basin boundary, where a trajectory’s final state is most sensitive to perturbations, does the RC occasionally predict the incorrect basin.
@@
Figure 4: In many scenarios, RC performance in unseen basins is comparable to performance in the training basin. We train an RC of $N_{r}=200$} nodes to forecast partially-observed states of the Duffing system ($x$ only) and plot the fraction of short test signals for which the RC predicts the correct basin of attraction as we vary the number of training initial conditions, $N_{train}$, and the half-widths of the ranges of the training and test initial conditions, $\Delta_{0}^{train}$} and $\Delta_{0}^{test}$}. (a) and (c): All training signals are from $B(\bm{A}^{-})$}. (b) and (d): Training signals are from both basins, $B(\bm{A}^{\pm})$}. In all cases, the training initial conditions are random, but the initial conditions of the test signals, belonging to both basins, form a $50\times 50$} uniform grid. Each test signal consists of $N_{test}=10$} consecutive data points from the true system, starting from the specified initial condition. For all grid points, we plot the mean fraction correct calculated over ten random realizations of the RC’s internal structure and of the training initial conditions. (a) and (b): $N_{train}=10$}. (c) and (d): $\Delta_{0}^{test}=10$}. Vertical lines mark the distance of the fixed points from the origin, $\bm{A}^{\pm}||$, and diagonal lines mark $\Delta_{0}^{test}=\Delta_{0}^{train}$}. Crosses and asterisks indicate the values of $N_{train}$, $\Delta_{0}^{train}$, and $\Delta_{0}^{test}$ that we use in Fig.˜3a,d,e and in Fig.˜3b-c, respectively